OpenStudy (anonymous):

Lastly, this question: 1/2x + 1/3y=6 & x-y=2 ( This is also solving systems by substitution, I have trouble with it!)

OpenStudy (anonymous):

Originally I got my answer as (6 2/5, 8 2/5) but apparently the answer is (8,6)! I need to understand the work for it.

OpenStudy (texaschic101):

x - y = 2 x = y + 2 1/2x + 1/3y = 6 -- lets get rid of the fractions by multiplying the entire equation by the LCD, which is 6. We will then get : 3x + 2y = 36 now we are gonna sub y + 2 in for x 3(y + 2) + 2y = 36 3y + 6 + 2y = 36 5y = 30 y = 6 x - y = 2 x - 6 = 2 x = 6 + 2 x = 8 any questions ? You can solve it by leaving things in fraction form, but I, myself, would rather not have the fractions.

OpenStudy (anonymous):

Thank you very much!

OpenStudy (texaschic101):

glad to help :) Just remember, if you have equations with fractions, you can multiply it by the LCD and you can get rid of the fractions

OpenStudy (anonymous):

Say for example I get the Regents test for Algebra. Will I get points off for leaving it as a fraction?

OpenStudy (texaschic101):

no..you should arrive at the same answer. Would you like me to do the problem with the fractions left in ?

OpenStudy (anonymous):

No thank you, because I realize that you are supposed to get rid of the fractions by multiplying the LCD and then you get (8,6)!

OpenStudy (anonymous):

But thank you though!

OpenStudy (texaschic101):

you will get the same answer if you leave it in fraction form